OpenStudy (anonymous):

1. Suppose that y varies inversely with x, and y = 6 when x = 8. What is an equation for the inverse variation? (1 point)

OpenStudy (anonymous):

Y = k/x 6=k/8 k=48 y=48/x

jimthompson5910 (jim_thompson5910):

y = k/x 6 = k/8 6*8 = k 48 = k k = 48 So the equation is \[\Large y = \frac{48}{x}\]